OpenStudy (anonymous):

What is the smallest positive integer less than 100 that has a remainder of 3 when divided by 7? How do you know this?

OpenStudy (anonymous):

10 10/7 = 1*7 + 3

OpenStudy (anonymous):

That's what I answered but the book has 3 and I'm not sure why

OpenStudy (anonymous):

ok, see if you try to divide 3 by seven you cannot this can be represented as 3 = 0*7 + 3

OpenStudy (anonymous):

ohhhh, ok! why didn't i get that *slaps head*

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

welcome ☺